OpenStudy (anonymous):

-[-(3x-5)+x] for x = -9 step by step please thanks

OpenStudy (chaise):

-[-(3x-5)+x] Simple substitute in your value of x=-9 -[-(3x-5)+-9] = -[-3x+5-9] = -[-3x-4] = [3x+4] When you see a negative outside a bracket assume you are multiplying every term in the bracket by -1.

OpenStudy (anonymous):

well what does it equal lol

OpenStudy (chaise):

3x+4. Sorry. Ignore my square brackets.

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

but wait you didnt substitute the 3x ?

OpenStudy (anonymous):

are you there lol

OpenStudy (chaise):

Oh, MY BAD. I'm sorry. = -[-((3x-9)-5)-9] = -[-((-27)-5)-9] = -[-(-32)] = -32